OpenStudy (anonymous):

Huilan is 9 years older than Thomas. The sum of their ages is 89. What is Thomas's age??

OpenStudy (anonymous):

is it 80

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

You can define the age of Thomas by T.

OpenStudy (anonymous):

Since Huilan is 9 years older than Thomas, you can represent his age by T + 9 to only have one variable in your equation.

OpenStudy (anonymous):

its 40 i got it

OpenStudy (anonymous):

Now, the sum of them two are 89, so we know both of their ages combined is equivalent to 89.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x = Thomas's age x+9 = Huilan's age Thomas + Huilan's age = 89 so x+x+9=89 2x+9=89 2x=80 x=40 Thomas's age X+9 or 49 is Hulians age

OpenStudy (anonymous):

(T+9) + T = 89 2T + 9 = 89 2T = 89 - 9 T = 80/2 T = 40